# Strange Quark Stars and Condensate Dark Stars in Bumblebee Gravity

### v1

@@ Abstract
In this paper, we investigate the properties of relativistic stars made of isotropic matter within the framework of the minimal Standard Model Extension, where a bumblebee field coupled to spacetime induces spontaneous Lorentz symmetry breaking. We adopt analytic equations-of-state describing either condensate dark stars or strange quark stars. We solve the structure equations numerically, and we compute the mass-to-radius relationships. The influence of the bumblebee parameter $l$ is examined in detail, and an upper bound is obtained using the massive pulsar PSR J0740+6620 and the strangely light HESS J1731-347 compact object.
@@ I Introduction
General Relativity and the Standard Model respectively describe gravity classically and particles/interactions quantum mechanically. Unifying these theories remains a fundamental challenge. While proposed quantum gravity theories offer potential solutions, direct experimental verification is currently unattainable due to the requisite Planck scale ($10^{19}$ GeV) energies. Nonetheless, subtle quantum gravity effects might manifest at lower energy scales, opening avenues for experimental investigation. One such potential signal is the breaking of Lorentz symmetry, rooted in special relativity, ensures that physical laws are consistent for all observers in inertial frames. This symmetry, encompassing rotational and boost transformations, is fundamental to general relativity and the standard model of particle physics . In curved spacetimes, local Lorentz symmetry is maintained due to the Lorentzian nature of the background. However, violating Lorentz invariance introduces directional or velocity dependencies in physical variables, altering the dynamics of particles and waves. In 1989, Kostelecký and Samuel introduced bumblebee gravity Kostelecky:1988zi , a model for spontaneous Lorentz violation. In this model, a bumblebee field with a vacuum expectation value (VEV) breaks Lorentz symmetry through the dynamics of a single vector field, $B_{\mu}$. This approach suggests that investigating Lorentz symmetry violation could provide insights into Planck-scale physics, combining aspects of both General Relativity and the Standard Model within the framework of effective field theories known as the Standard Model Extension (SME), which offers a framework that includes all possible coefficients for Lorentz/CPT violation, with its gravitational sector defined on a Riemann-Cartan manifold, treating torsion and the metric as dynamic geometric quantities. While the SME allows for non-Riemannian terms, research has primarily focused on the metric approach, using the metric as the main dynamic field. There are certain effects of Lorentz violation within the gravitational sector have been explored in Kostelecky:2000mm ; Kostelecky:2001mb ; Bluhm:2004ep ; Kostelecky:1989jw ; AraujoFilho:2024ykw ; Filho:2022yrk ; Delhom:2022xfo ; Delhom:2019wcm ; Lambiase:2023zeo ; Gullu:2020qzu ; Li:2020dln ; Maluf:2020kgf ; Maluf:2014dpa ; Casana:2017jkc ; Bertolami:2005bh ; Capozziello:2023tbo ; Capozziello:2023rfv .
@@
New discoveries of super-dense objects are pushing the boundaries of our understanding in astrophysics. While countless compact objects have been found, the exact laws governing their behavior remain a mystery. Recent observations have challenged the traditional theories, particularly general relativity and the standard model for neutron stars. One key example is the secondary object in the GW190814 merger Zhang:2020jmb ; Zhang:2021iah ; Gammon:2024gij . Some scientists believe it might be a quark star, with its properties described by a single parameter ($\lambda$) within the framework of general relativity. This theory aligns with recent research suggesting non-strange quark matter as the most stable form of baryonic matter under extreme conditions Holdom:2017gdc . Promising findings have also been obtained by similar investigations using the quark matter phase or using other models for matter behaviour (equation of state) Casana:2017jkc ; Bertolami:2005bh ; Capozziello:2023tbo ; Capozziello:2023rfv ; Gammon:2023uss .
Strange stars: Strange quark stars are hypothetical compact objects proposed as an alternative to neutron stars Madsen:1998uh ; Yue:2006it ; Leahy:2007we ; Arapoglu:2010rz ; Eksi:2015red ; Azri:2021pxu ; Baym:2017whm . Composed of quark matter, which is theorized to be the absolutely stable ground state of hadrons Witten:1984rs ; Farhi:1984qu , these stars could explain the puzzling nature of super-luminous supernovae Ofek:2006vt ; Ouyed:2009dr . Unlike ordinary supernovae, these exceptionally bright explosions are 100 times more luminous and occur in approximately one out of every thousand cases.
@@
While strange quark stars remain theoretical at present, their existence cannot be definitively excluded. Some observed compact objects exhibit anomalous properties, such as unusually small radii, that standard neutron star models cannot fully explain Henderson:2007gu ; Li:2011zzn ; Maxim . These discrepancies have fueled speculation about the potential presence of strange matter in these objects. While some researchers propose that strange matter could constitute the core of hybrid neutron stars Benic:2014jia ; Yazdizadeh:2019ymr ; EslamPanah:2018rfe , others argue that such stars would be virtually indistinguishable from ordinary neutron stars Jaikumar:2005ne .
@@
Boson stars: The standard cosmological model posits that dark matter consists of weakly interacting massive particles (WIMPs). While this assumption aligns well with large-scale cosmic observations (scales of megaparsecs and larger) ($\geq Mpc$), it faces challenges at smaller galactic scales. Discrepancies such as the core-cusp problem, the diversity problem, the missing satellites problem, and the too-big-to-fail problem arise Tulin:2017ara . Self-interacting dark matter has been proposed as a potential solution Spergel:1999mh ; Dave:2000ar , as particle collisions within this model can mitigate the formation of sharp density cusps. Furthermore, if dark matter is composed of ultralight scalar particles with a mass below one electronvolt $m\leq eV$ and a weak repulsive self-interaction, it could form a Bose-Einstein condensate (BEC) with long-range correlations. This BEC scenario has been proposed as a potential explanation for the aforementioned galactic scale discrepancies Tkachev:1986tr ; Goodman:2000tg ; Peebles:2000yy .
@@
Our current understanding of ultra-dense stars, particularly those exceeding the limitations of traditional neutron stars, remains a fascinating mystery in astrophysics. This work explores this field by examining the characteristics of isotropically-composed stars.
Our key objective is to explore the impact of a phenomenon known as spontaneous Lorentz symmetry breaking within these exotic stars. This breaking is induced by a hypothetical field, the bumblebee field, coupled to spacetime in the framework of the minimal Standard Model Extension.
@@ II Relativistic stars in General Relativity: Hydrostatic equilibrium and structure equations
Due to their intense gravitational fields, the structure and dynamics of neutron stars (NSs) are governed by Einstein’s equations of General Relativity (GR) Einstein:1915ca , given by:
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi T_{\mu\nu},$ (1)
@@
where $g_{\mu\nu}$ is the metric tensor, $P$ is the pressure, $\mathcal{E}$ is the energy density, and $u_{\mu}$ is the four-velocity. For static, spherically symmetric stars, the line element in Schwarzschild-like coordinates $\{t,r,\theta,\phi\}$ is described as:
$ds^{2}=-e^{\nu(r)}dt^{2}+e^{\lambda(r)}dr^{2}+r^{2}(d\theta^{2}+sin^{2}\theta d\phi^{2}),$ (3)
@@
Combined with the given EoS—i.e., $P(\rho)$ of the matter—, TOV equations can be solved with the initial conditions at the center of the star, $m(r=0)$ = 0 and $\rho(r=0)$ = $\rho_{c}$, where $\rho_{c}$ is the central energy density. The stellar mass and radius are determined using the matching conditions at the surface of the star upon comparison to the Schwarzschild exterior vacuum solution Schwarzschild:1916uq
$ds^{2}=-(1-2M/r)\>dt^{2}+\frac{1}{1-2M/r}\>dr^{2}+r^{2}d\Omega^{2}.$ (8)
@@ III Bumblebee gravity: Modified TOV equations
In this section, we briefly review the bumblebee gravity, a theory that expands on General Relativity (GR). Inspired by the works of Kostelecky and collaborators Kostelecky:1988zi , bumblebee gravity introduces a twist: it breaks a fundamental symmetry (Lorentz symmetry) within the realm of gravity. This twist manifests as a special value (nontrivial vacuum expectation value) that influences how other fields behave around a mysterious ”bumblebee field”. Interestingly, even with these extra interactions, bumblebee gravity preserves the geometrical framework and basic laws established by GR in curved spaces.
Among various models capable of breaking Lorentz symmetry, one of the simplest approaches involves a vector field $B^{\mu}$, known as the bumblebee field, within a torsion-free spacetime. This can be expressed as follows Casana:2017jkc :
$\begin{split}\mathcal{S}=&\int\mathrm{d}^{4}x\sqrt{\left|g\right|}\left(\frac{1}{16\pi}\left(\mathcal{R}+\xi B^{\mu}B^{\nu}\mathcal{R}_{\mu\nu}\right)-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V\left(B^{\mu}B_{\mu}\pm b^{2}\right)+\mathcal{L}_{M}\right),\end{split}$ (9)
which encapsulates the interaction between the bumblebee vector field ($B_{\mu}$) and gravity, represented by the Ricci tensor ($R_{\mu\nu}$). The strength of this interaction is modulated by a coupling constant ($\xi$).
The bumblebee field $B_{\mu\nu}\equiv\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$ also interacts with matter, as outlined by the matter Lagrangian density ($L_{m}$). The potential function $V(B_{\mu}B_{\nu}\pm b^{2})$, where $b^{2}$ is a positive constant, dictates the behavior of the bumblebee field. Notably, this potential enables the field to achieve a non-zero vacuum expectation value, specifically $B_{\mu}=b_{\mu}=(0,b_{r}(r),0,0)$, where $b_{r}(r)$ is a function of the radial coordinate.
To break a specific symmetry, such as $U(1)$, the potential must reach a minimum at $V=0$, with its derivative $V^{\prime}=0$ also being zero at that point. This condition allows the bumblebee field to acquire a constant, non-zero vacuum expectation value, denoted by $b_{\mu}$, which satisfies ${B^{\mu}}\equiv b^{\mu}$ with $b^{\mu}b_{\mu}=\mp b^{2}\equiv$ constant (note that $b^{2}$ is a real positive constant).
By varying the action, as given in Eq. (9), we derive the equations governing both the gravitational field and the dynamics of the bumblebee field, including the corresponding equations of motion.
@@
$\nabla_{\mu}B^{\mu\nu}=\mathcal{J}^{B}_{\nu}+\mathcal{J}^{M}_{\nu}.$ (11)
Within the bumblebee gravity framework, two key terms come into play: the source term for the BF (represented by $\mathcal{J}^{M}_{\nu}$) and the current due to the BF’s self-interaction (denoted by $\mathcal{J}^{B}_{\nu}=2V^{\prime}B^{\nu}-\varrho B_{\mu}\mathcal{R}^{\mu\nu}/8\pi$). Note that the prime symbol ($\prime$) denotes differentiation with respect to the potential argument, specifically $V^{\prime}\equiv\partial V(y)/\partial y$, where $y=B^{\mu}B_{\nu}\pm b^{2}$.
Additionally, $T^{B}_{\mu\nu}$ represents the energy-momentum tensor associated with the bumblebee field, which can be expressed as
@@
Remember that $T_{\mu\nu}^{M}$ represents the energy and momentum distribution of matter, and we have a constant value denoted by $b^{\mu}b_{\mu}=\mp b^{2}\equiv\text{constant}$ related to the bumblebee field (where $b_{\mu}$ is the VEV and b is its magnitude). These factors influence the radial component $b_{r}(r)$ of the BF when it reaches its minimum energy state (VEV). Interestingly, research by Casana et al. Casana:2017jkc shows that this radial component can be expressed as $b_{r}(r)=|b|e^{\zeta(r)}$.
@@
$\begin{split}\text{g}(r)=1-\frac{2\,\mathrm{M}(r)}{r},\end{split}$ (14)
in which the collective mass function denoted by $\mathrm{M}(r)$ can be viewed as the combination of the BH mass.
Thus, for an interacting system of this nature, the energy-momentum tensor associated with the spacetime metric can be represented as $T^{\mu}_{\,\,\nu}=\text{diag}\left[-\rho(r),P_{r}(r),P_{\theta}(r),P_{\phi}(r)\right]$. Consequently, these considerations facilitate the derivation of the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$ as
$8\pi\rho(r)$ $=\frac{1-q^{2}+2q^{2}\text{M}^{\prime}(r)}{q^{2}r^{2}},$ (15a)
$8\pi P_{r}(r)$ $=\frac{1}{r^{2}}\left(1-\frac{1}{q^{2}}\right)-\frac{2\text{M}(r)}{r^{3}}+\frac{2}{r}\left(1-\frac{2\text{M}(r)}{r}\right)\chi^{\prime}(r),$ (15b)
$8\pi P_{\theta}(r)=8\pi P_{\varphi}(r)$ $=\left(\frac{1}{r}+\chi^{\prime}(r)\right)\left(\frac{\text{M}(r)}{r^{2}}-\frac{\text{M}^{\prime}(r)}{r}+\left(1-\frac{2\text{M}(r)}{r}\right)\chi^{\prime}(r)\right)$
$+\left(1-\frac{2\text{M}(r)}{r}\right)\chi^{\prime\prime}(r).$ (15c)
Within the bumblebee gravity framework, the parameter called $q^{2}=1/(l+1)$ plays a crucial role. It depends on a constant known as parameter of Lorentz symmetry breaking $l$, which itself is linked to the coupling strength $\xi$ of the BF and a constant related to the BF $b\texttwosuperior$ ² as with $l=\xi b^{2}$. The prime symbol $(^{\prime})$ denotes differentiation with respect to the radial position (r). Interestingly, when we take the limit as $l$ approaches zero, $q\texttwosuperior$ ² approaches $1$, has a significant implication. In this specific limit ($l\rightarrow 0,q^{2}\rightarrow 1$), the Einstein field equations (refer to Eq. (15) for reference) simplify back to their standard form.
Now, let us bring in another piece of the puzzle: the equation of state $P_{r}(r)=\omega\rho(r)$, which relates pseudo-pressure ($P_{r}(r)$) and energy density ($\rho(r)$) with a constant factor ($\omega$). By combining this equation of state with the modified Einstein field equations we obtained earlier (Eq. (15a)), and considering a well-known law in physics (conservation of energy-momentum) $T_{\mu;\nu}^{\nu}=0$, we can arrive at a modified version of the TOV equations. These modified TOV equations are crucial for understanding how bumblebee gravity affects the structure and stability of stars and other self-gravitating objects.
$\frac{d\chi(r)}{dr}$ $=\,\frac{\text{M}(r)-\left(\frac{1}{2}-\frac{1}{2q^{2}}\right)r+4\pi r^{3}P_{r}(r)}{r(r-2\text{M}(r))},$ (16a)
$\frac{dP_{r}(r)}{dr}$ $=-\left(\rho(r)+P_{r}(r)\right)\frac{d\chi(r)}{dr},$ (16b)
Hence, in the exterior (vacuum) region where $T_{\mu\nu}=0$, the solution simplifies to the bumblebee black hole (BH) solution Casana:2017jkc
$ds^{2}=-(1-2M/r)\>dt^{2}+\frac{l+1}{1-2M/r}\>dr^{2}+r^{2}d\Omega^{2},$ (17)
or, after a coordinate transformation and a mass redefinition as follows
@@
$ds^{2}=-(1-2M/r)\>dt^{2}+\frac{1}{1-2M/r}\>dr^{2}+q^{2}r^{2}d\Omega^{2}.$ (19)
When studying a star in bumblebee gravity, we need to ensure our solutions for the different regions (interior and exterior) seamlessly match at the star’s surface, located at a specific radius ($R$). To achieve this perfect match, we impose three matching conditions when $q\neq 1$:

### v2

@@ Abstract
In this paper, we investigate the properties of relativistic stars made of isotropic matter within the framework of the minimal Standard Model Extension, where a Bumblebee field (BF) coupled to spacetime induces spontaneous Lorentz symmetry breaking. We adopt analytic equations-of-state describing either condensate dark stars or strange quark stars. We solve the structure equations numerically, and we compute the mass-to-radius relationships. The influence of the Bumblebee parameter $\mathbf{l}$ is examined in detail, and an upper bound is obtained using the massive pulsar (PSR) J0740+6620 and the strangely light High Energy Stereoscopic System (HESS) J1731-347 compact object.
@@ I Introduction
General Relativity (GR) and the Standard Model (SM) of Particle Physics, respectively, describe gravity classically and particles/interactions quantum mechanically. Unifying these theories remains a fundamental challenge. While proposed quantum gravity theories offer potential solutions, direct experimental verification is currently unattainable due to the requisite Planck scale ($10^{19}$ GeV) energies. Nonetheless, subtle quantum gravity effects might manifest at lower energy scales, opening avenues for experimental investigation. One such potential signal is the breaking of Lorentz symmetry, rooted in special relativity, ensures that physical laws are consistent for all observers in inertial frames. This symmetry, encompassing rotational and boost transformations, is fundamental to general relativity and the standard model of particle physics. In curved space-times, local Lorentz symmetry is maintained due to the Lorentzian nature of the background. However, violating Lorentz invariance introduces directional or velocity dependencies in physical variables, altering the dynamics of particles and waves. In 1989, Kostelecký and Samuel introduced Bumblebee gravity Kostelecky:1988zi , a model for spontaneous Lorentz violation. In this model, a BF with a vacuum expectation value (VEV) breaks Lorentz symmetry through the dynamics of a single vector field, $B_{\mu}$. This approach suggests that investigating Lorentz symmetry violation could provide insights into Planck-scale physics, combining aspects of both General Relativity and the Standard Model within the framework of effective field theories known as the Standard Model Extension (SME), which offers a framework that includes all possible coefficients for Lorentz/CPT violation, with its gravitational sector defined on a Riemann-Cartan manifold, treating torsion and the metric as dynamic geometric quantities. While the SME allows for non-Riemannian terms, research has primarily focused on the metric approach, using the metric as the main dynamic field. There are certain effects of Lorentz violation within the gravitational sector have been explored in Kostelecky:2000mm ; Kostelecky:2001mb ; Bluhm:2004ep ; Kostelecky:1989jw ; AraujoFilho:2024ykw ; Filho:2022yrk ; Delhom:2022xfo ; Delhom:2019wcm ; Lambiase:2023zeo ; Gullu:2020qzu ; Li:2020dln ; Maluf:2020kgf ; Maluf:2014dpa ; Casana:2017jkc ; Bertolami:2005bh ; Capozziello:2023tbo ; Capozziello:2023rfv .
@@
New discoveries of super-dense objects are pushing the boundaries of our understanding in Astrophysics. While countless compact objects have been found, the exact laws governing their behavior remain a mystery. Recent observations have challenged the traditional theories, particularly general relativity and the standard model for neutron stars (NSs). One key example is the secondary object in the GW190814 merger Zhang:2020jmb ; Zhang:2021iah ; Gammon:2024gij , which has been proposed as a potential quark star. Its properties have been analyzed in the framework of general relativity, with models suggesting that they may be described by a single parameter ($\lambda$). This theory aligns with recent research suggesting non-strange quark matter as the most stable form of baryonic matter under extreme conditions Holdom:2017gdc . Promising findings have also been obtained by similar investigations using the quark matter phase or using other models for matter behavior (equation-of-state) Casana:2017jkc ; Bertolami:2005bh ; Capozziello:2023tbo ; Capozziello:2023rfv ; Gammon:2023uss .
@@
Strange quark stars are hypothetical compact objects proposed as an alternative to neutron stars Madsen:1998uh ; Yue:2006it ; Leahy:2007we ; Arapoglu:2010rz ; Eksi:2015red ; Azri:2021pxu ; Baym:2017whm ; Pradhan:2024zzo ; Bhar:2023yrf ; Panotopoulos:2024iag ; Rincon:2023ens ; Rahaman:2011cw ; Sedaghat:2022fue ; Sedaghat:2024bnj ; Deb:2017rhd ; Bhar:2014tqa . Composed of quark matter, which is theorized to be the absolutely stable ground state of hadrons Witten:1984rs ; Farhi:1984qu , these stars could explain the puzzling nature of super-luminous supernovae Ofek:2006vt ; Ouyed:2009dr . Unlike ordinary supernovae, these exceptionally bright explosions are 100 times more luminous and occur in approximately one out of every thousand cases. According to the standard terminology, in the literature the names ”quark stars”, ”strange stars” and ”strange quark stars” are used to refer to the same object. Strange quark stars are made of u, d and s quarks, as the other three flavors are too heavy. Their mass is much larger than the chemical potentials involved, and so they are not relevant. While quark stars remain theoretical at present, their existence cannot be definitively excluded. Some observed compact objects exhibit anomalous properties, such as unusually small radii, that standard neutron star models cannot fully explain Henderson:2007gu ; Li:2011zzn ; Maxim . These discrepancies have fueled speculation about the potential presence of strange matter in these objects. While some researchers propose that strange matter could constitute the core of hybrid neutron stars Benic:2014jia ; Yazdizadeh:2019ymr ; EslamPanah:2018rfe , others argue that such stars would be virtually indistinguishable from ordinary neutron stars Jaikumar:2005ne .
@@
The standard cosmological model posits that dark matter consists of weakly interacting massive particles (WIMPs). While this assumption aligns well with large-scale cosmic observations (scales of megaparsecs and larger) ($\geq Mpc$), it faces challenges at smaller galactic scales. A few discrepancies between observations and model predictions, such as the core-cusp problem, the diversity problem, the missing satellites problem, and the too-big-to-fail problem, arise Tulin:2017ara . Self-interacting dark matter has been proposed as a potential solution Spergel:1999mh ; Dave:2000ar , as particle collisions within this model can mitigate the formation of sharp density cusps. Furthermore, if dark matter is composed of ultralight scalar particles with a mass below one electronvolt $m\leq eV$ and a weak repulsive self-interaction, it could form a Bose-Einstein condensate (BEC) harko1 with long-range correlations. This BEC scenario has been proposed as a potential explanation for the dark matter crisis on small (i.e. galactic) scales Tkachev:1986tr ; Goodman:2000tg ; Peebles:2000yy ; harko2 .
@@
Our current understanding of ultra-dense stars, particularly those exceeding the limitations of traditional neutron stars, remains a fascinating mystery in Astrophysics. Since compact objects are characterized by high matter densities and strong gravitational fields, they are excellent cosmic laboratories to constrain gravitational theories other than GR. Since the astrophysical implications of the Bumblebee gravity should also be investigated, in this work we propose to study the structural properties of compact stars made of isotropic matter, and in particular those of quark stars and self-gravitating spheres made of dark matter condensate harko3 ; harko4 .
Our key objective is to explore the impact of a phenomenon known as spontaneous Lorentz symmetry breaking within these exotic stars. This breaking is induced by a hypothetical field, the BF, coupled to spacetime in the framework of the minimal SME.
@@ II Relativistic stars in General Relativity: Hydrostatic equilibrium and structure equations
Due to their intense gravitational fields, the structure and dynamics of NSs are governed by Einstein’s equations of GR Einstein:1915ca , given by:
$G_{\mu\nu}=\mathcal{R}_{\mu\nu}-\frac{1}{2}\mathcal{R}g_{\mu\nu}=8\pi T_{\mu\nu},$ (1)
@@
where $g_{\mu\nu}$ is the metric tensor, $P$ is the pressure, $\rho$ is the energy density, and $u_{\mu}$ is the four-velocity. For static, spherically symmetric stars, the line element in Schwarzschild-like coordinates $\{t,r,\theta,\phi\}$ is described as:
$ds^{2}=-e^{\nu(r)}dt^{2}+e^{\lambda(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (3)
@@
Combined with the given the equation of state (EoS)—i.e., $P(\rho)$ of the matter—, TOV equations can be solved with the initial conditions at the center of the star, $m(r=0)$ = 0 and $\rho(r=0)$ = $\rho_{c}$, where $\rho_{c}$ is the central energy density. The stellar mass and radius are determined using the matching conditions at the surface of the star upon comparison to the Schwarzschild exterior solution in bumblebee gravity Schwarzschild:1916uq
$ds^{2}=-(1-2M/r)\>dt^{2}+(1+\mathbf{l})\frac{1}{1-2M/r}\>dr^{2}+r^{2}d\Omega^{2}.$ (8)
@@ III Bumblebee gravity: Modified TOV equations
In this section, we briefly review the bumblebee gravity, a theory that expands on GR. Inspired by the works of Kostelecky and collaborators Kostelecky:1988zi , bumblebee gravity introduces a twist: it breaks a fundamental symmetry (Lorentz symmetry) within the realm of gravity. This twist manifests as a special value (nontrivial vacuum expectation value) that influences how other fields behave around a mysterious ”bumblebee field”. Interestingly, even with these extra interactions, bumblebee gravity preserves the geometrical framework and basic laws established by GR in curved spaces.
Among various models capable of breaking Lorentz symmetry, one of the simplest approaches involves a vector field $B^{\mu}$, known as the BF, within a torsion-free spacetime. This can be expressed as follows Casana:2017jkc :
$\begin{split}\mathcal{S}=&\int\mathrm{d}^{4}x\sqrt{\left|g\right|}\left(\frac{1}{16\pi}\left(\mathcal{R}+\xi B^{\mu}B^{\nu}\mathcal{R}_{\mu\nu}\right)-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V\left(B^{\mu}B_{\mu}\pm b^{2}\right)+\mathcal{L}_{M}[g]\right),\end{split}$ (9)
which encapsulates the interaction between the bumblebee vector field ($B_{\mu}$) and gravity, represented by the Ricci tensor ($R_{\mu\nu}$). The strength of this interaction is modulated by a coupling constant ($\xi$). Notice that in this work we assume that the Lagrangian of matter content only depends on the metric tensor $g_{\mu\nu}$, and it is $B_{\mu}$ independent.
The BF $B_{\mu\nu}\equiv\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$ does not coupled with matter, as outlined by the matter Lagrangian density ($L_{m}$) which contains a perfect fluid. The potential function $V(B_{\mu}B_{\nu}\pm b^{2})$, where $b^{2}$ is a positive constant, dictates the behavior of the BF. Notably, this potential enables the field to achieve a non-zero vacuum expectation value, specifically $B_{\mu}=b_{\mu}=(0,b_{r}(r),0,0)$, where $b_{r}(r)$ is a function of the radial coordinate.
To break a specific symmetry, such as $U(1)$, the potential must reach a minimum at $V=0$, with its derivative $V^{\prime}=0$ also being zero at that point. This condition allows the BF to acquire a constant, non-zero vacuum expectation value, denoted by $b_{\mu}$, which satisfies ${B^{\mu}}\equiv b^{\mu}$ with $b^{\mu}b_{\mu}=\mp b^{2}\equiv$ constant (note that $b^{2}$ is a real positive constant).
Varying the action, as given in Eq. (9), we derive the equations governing both the gravitational field and the dynamics of the BF, including the corresponding equations of motion.
@@
$\nabla_{\mu}B^{\mu\nu}=\mathcal{J}_{B}^{\nu}+\mathcal{J}_{M}^{\nu}.$ (11)
In the Bumblebee gravity framework, two essential terms are: the matter source term associated with the Bumblebee field, represented by $\mathcal{J}^{M}_{\nu}$, and the self-interaction current of the Bumblebee field (BF), denoted by $\mathcal{J}^{B}_{\nu}=2V^{\prime}B^{\nu}-(\varrho/8\pi)B_{\mu}\mathcal{R}^{\mu\nu}$.
Given that the field strength tensor vanishes, $B_{\mu\nu}=0$, Eq. 11 yields the relation $\mathcal{J}^{\mu}_{B}=-\mathcal{J}^{\mu}_{M}$. We assume that $|\mathcal{J}^{\mu}_{B}|\approx 0$, indicating negligible interaction. Therefore, it is reasonable to assume the absence of coupling between the matter sector and the Bumblebee field. This assumption is further supported by the fact that the matter Lagrangian $\mathcal{L}_{m}$ in our model describes a perfect fluid. This scenario aligns with the approach previously considered in Ref. Neves:2024ggn .
The prime symbol () denotes differentiation with respect to the potential argument, specifically $V^{\prime}\equiv\partial V(y)/\partial y$, where $y=B^{\mu}B_{\mu}\pm b^{2}$.
Additionally, $T^{B}_{\mu\nu}$ represents the energy-momentum tensor associated with the BF, which can be expressed as
@@
Recall that $T_{\mu\nu}^{M}$ represents the energy-momentum distribution of matter, while $\varrho$ is the real coupling constant (with mass dimension $-1$) that governs the non-minimal interaction between gravity and the bumblebee field. Additionally, we introduce a constant value denoted by $b^{\mu}b_{\mu}=\mp b^{2}\equiv\text{constant}$ related to the BF (where $b_{\mu}$ is the VEV and b is its magnitude). These factors influence the radial component $b_{r}(r)$ of the BF when it reaches its minimum energy state (VEV). Interestingly, research by Casana et al. Casana:2017jkc shows that this radial component can be expressed as $b_{r}(r)=|b|e^{\zeta(r)}$.
@@
$\begin{split}g(r)=1-\frac{2\,M(r)}{r},\end{split}$ (14)
in which the collective mass function denoted by $\mathrm{M}(r)$ can be viewed as the combination of the black hole (BH) mass.
Thus, for an interacting system of this nature, the energy-momentum tensor associated with the spacetime metric can be represented as a perfect fluid $T^{\mu}_{\,\,\nu}=\text{diag}\left[-\rho(r),P_{r}(r),P_{\theta}(r),P_{\phi}(r)\right]$. Consequently, these considerations facilitate the derivation of the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$ as
$8\pi\rho(r)$ $=\frac{e^{-2\zeta(r)}}{r^{2}}\left[e^{2\zeta(r)}-(1+\mathbf{l})\left(1-2r\zeta(r)^{\prime}\right)\right]=\kappa\rho$ (15a)
$8\pi P_{r}(r)$ $=\frac{e^{-2\zeta(r)}}{r^{2}}\left[(1+\mathbf{l})\left(1+2r\chi(r)^{\prime}\right)-e^{2\zeta(r)}-\ell r^{2}\left(\chi(r)^{\prime\prime}+\chi(r)^{\prime 2}-\chi(r)^{\prime}\zeta(r)^{\prime}-\frac{2}{r}\zeta(r)^{\prime}\right)\right]$ (15b)
$8\pi P_{\theta}(r)=8\pi P_{\varphi}(r)$ $=(1+\mathbf{l})e^{-2\zeta(r)}\left[\chi(r)^{\prime\prime}+\chi(r)^{\prime 2}-\chi(r)^{\prime}\zeta(r)^{\prime}+\frac{1}{r}\left(\chi(r)^{\prime}-\zeta(r)^{\prime}\right)\right].$ (15c)
Within the bumblebee gravity framework, parameter of Lorentz symmetry breaking is $\mathbf{l}$, which itself is linked to the coupling strength $\xi$ of the BF and a constant related to the BF $b\texttwosuperior$ ² as with $\mathbf{l}=\xi b^{2}$. The prime symbol $(^{\prime})$ denotes differentiation with respect to the radial position (r). Interestingly, when we take the limit as $\mathbf{l}$ approaches zero, has a significant implication. In this specific limit ($\mathbf{l}\rightarrow 0$ ), the Einstein field equations (refer to Eq. (15) for reference) simplify back to their standard form.
Now, let us bring in another piece of the puzzle: the equation of state $P_{r}(r)=\omega\rho(r)$, which relates pseudo-pressure ($P_{r}(r)$) and energy density ($\rho(r)$) with a constant factor ($\omega$). By combining this equation of state with the modified Einstein field equations we obtained earlier (Eq. (15c)), and considering a well-known law in physics (conservation of energy-momentum) $T_{\mu;\nu}^{\nu}=0$, we can arrive at a modified version of the TOV equations.These modified TOV equations are crucial for understanding how bumblebee gravity affects the structure and stability of stars and other self-gravitating objects Neves:2024ggn :
$\frac{dM(r)}{dr}$ $=4\pi r^{2}\,\rho(r),$ (16a)
[…]
and
$\chi^{\prime}(r)=\frac{(1+2\mathbf{l})8\pi r^{3}P_{r}(r)+(2+3\mathbf{l})M(r)-\mathbf{l}rM^{\prime}(r)}{(2+3\mathbf{l})r(r-2M(r))}.$ (17)
Hence, in the exterior (vacuum) region where $T_{\mu\nu}=0$, the geometry simplifies to the bumblebee BH solution Casana:2017jkc
$ds^{2}=-(1-2M/r)\>dt^{2}+\frac{\mathbf{l}+1}{1-2M/r}\>dr^{2}+r^{2}d\Omega^{2},$ (18)
or, after a coordinate transformation and a stellar mass redefinition as follows
@@
$ds^{2}=-(1-2M/r)\>dt^{2}+\frac{1}{1-2M/r}\>dr^{2}+q^{2}\>r^{2}d\Omega^{2},$ (20)
where $q^{2}=1/(\mathbf{l}+1)$, and where now the modification due to the Bumblebee gravity is relocated to the angular part of the line element. In the discussion to follow, however, we shall employ the more standard coordinate system, namely the previous form of the metric tensor without the coordinate transformation and the mass redefinition.
This result indicating that a background field carrying Lorentz-violating effects also slightly deforms spacetime. Despite these effects being significantly small, with a sensitivity at the $10^{-13}$ level, spacetime is described as approximately flat. This smallness has facilitated the computation of upper bounds on a parameter $\mathbf{l}$, referenced in Casana:2017jkc .
When studying a star in Bumblebee gravity, we need to ensure our solutions for the different regions (interior and exterior) seamlessly match at the star’s surface, located at a specific radius ($R$). To achieve this perfect match, we impose three matching conditions as follows: